[10] 6. Find the general solutions to the equations (a) 2x+y + 5xy + y = 0, (6) Y – 2y + 5y = 3, YIV – 2k²y + k*y = 0, (c
Answer #1

solution aj Given differential equr is 2x²y + 5*4l +y=0 which is -(1) a homogeneous linear D.E with variable coefficient the3 for a2+bx+C =0 xa -b I 562_fac 2X) 2a - 2 ITC 2 Hence, general solution Auxiliary equation is - m² 2m +5=0 m- 27V €2)24x) X